Required information [The following information applies to the questions displayed below.] The general ledger of Jackrabbit Rentals at January 1, 2021, includes the following account balances: Accounts Debits Credits Cash $ 46,500 Accounts Receivable 30,700 Land 115,800 Accounts Payable 15,800 Notes Payable (due in 2 years) 35,000 Common Stock 105,000 Retained Earnings 37,200 Totals $ 193,000 $ 193,000 The following is a summary of the transactions for the year: 1. January 12 Provide services to customers on account, $67,400. 2. February 25 Provide services to customers for cash, $77,800. 3. March 19 Collect on accounts receivable, $46,200. 4. April 30 Issue shares of common stock in exchange for $35,000 cash. 5. June 16 Purchase supplies on account, $13,100. 6. July 7 Pay on accounts payable, $11,800. 7. September 30 Pay salaries for employee work in the current year, $69,200. 8. November 22 Pay advertising for the current year, $23,000. 9. December 30 Pay $3,400 cash dividends to stockholders. The following information is available for the adjusting entries. Accrued interest on the notes payable at year-end amounted to $3,000 and will be paid January 1, 2022. Accrued salaries at year-end amounted to $2,000 and will be paid on January 5, 2022. Supplies remaining on hand at the end of the year equal $2,800.

8-a. Prepare an income statement for the year ended December 31, 2021.

8-b. Prepare a classified balance sheet for the year ended December 31, 2021.
Answer #1
8-a.) Income Statement
For the year ended December 31,2021 Amount $
Revenue
Service Revenue (77,800 + 67,400 )     145,200
Expenses
Salaries expense         69,200
Advertising Expense         23,000
Supplies Expense ( 13,100 - 2,800 )         10,300
Interest Expense           3,000
Total Expense     105,500
Net Income        39,700
8-b.) Classified Balance Sheet
For the year ended December 31,2021 Amount $
Assets
Current Assets
Cash         98,100 =46500+77800+46200+35000-11800-69200-23000-3400
Accounts Receivable         51,900 =30700+67400-46200
Supplies           2,800
Non-Current Assets
Land     115,800
Total Assets     268,600
Liabilities & Stockholder's Equity
Non-Current Liabilities
Notes Payable         35,000
Current Liabilities
Accounts Payable         17,100 =15800+13100-11800
Interest Payable           3,000
Stockholder's Equity
Common Stock     140,000 =35000+105000
Retained earnings         73,500 =37200-3400+39700
Total Liabilities & Stockholder's Equity     268,600